The potential of the cell Bi |Bi2S3(s)|Bi2S3(aq)|Bi is 0.96 V at 25°C. Calculate (a) the solubility product of Bi2S3 and (b) its solubility


1
Expert's answer
2021-11-08T08:11:59-0500

Contemporary Structures Creating a contemporary structure involves reconfiguring the firm’s value chain to… • Gain cost savings • Gain specialization benefits • Improve integration and coordination Margin  Support Activities: o Human Resources o Accounting and infrastructure o Procurement and technology  Primary Activities: o In-bound Logistics o Operations or Manufacturing o Out-bound Logistics o Marketing and Sales o Service and Support Contemporary structures also called networked or hollow structures There are a variety of contemporary organizational structures. Many of them do not have common labels or names. This is, in part, because many of them have their essence in the configuration of organic compounds.

= 0.96×298= 286.08/1000

= 2.8608 x 10-2 moles per liter